{Exercise 2.3} A questionnaire provides 58 Yes, 42 No, and 20 No-Opinion answers. In the construction of a pie chart, how many degrees would be in the section of the pie showing the Yes answers? degrees How many degrees would be in the section of the pie showing the No answers? degrees Construct a pie chart. Then, enter the percentages of the circle occupied by each response (to 1 decimal). Yes % No % No Opinion %

  1. In the construction of a pie chart, how many degrees would be in the section of the pie showing the Yes answers?
      degrees
  2. How many degrees would be in the section of the pie showing the No answers?
      degrees
  3. Construct a pie chart. Then, enter the percentages of the circle occupied by each response (to 1 decimal).
    Yes   %
    No   %
    No Opinion   %
Answer #1

a)

Degree of Yes answers = 58/120 * 360= 174 degrees

b)

Degree of No answers = 42/120 * 360 = 126 degrees

c)

The pie chart is given below:

Question? Result %
Yes 58 48.3%
No 42 35.0%
No Opinion 20 16.7%

Please let me know in comments in case anything is unclear. Will reply ASAP. Do upvote if satisfied!